For CuCO3, Ksp is 2.5 x 10-10 M2. For CuCl42-, Kf is 5.0 x 105 M-4. What is the value of the equilibrium constant for the sum of the two reactions shown below?

CuCO3 ↔ Cu2+ + CO32- Ksp = 2.5 x 10-10 M2

Cu2+ + 4 Cl- ↔ CuCl42- Kf = 5.0 x 105 M-4

CuCO3 + 4 Cl- ↔ CuCl42- + CO32- K = ?

K = ___ M-2

Based on the previous problem, what is the maximum solubility of CuCO3 in a 0.650 M NaCl solution?

S = ___ M
Answer #1

i) CuCO3(s) <------> Cu2+(aq) + CO32-(aq)   

Ksp =[Cu2+][CO32-] = 2.5×10-10M2

Cu2+(aq) + 4Cl-(aq) <--------> CuCl4-(aq)

Kf = [CuCl4-]/[Cu2+][Cl-]4 = 5.0×105M-4

adding two equation

CuCO3(s) + 4Cl-(aq) <---------> CuCl4-(aq) + CO32-(aq)

K = [CuCl4-][CO32-]/[Cl-]4

K = Ksp × Kf

K = 2.5×10-10M2× 5.0×105M-4

K = 1.25 × 10-4M-2

ii) CuCO3(s) + 4Cl- (aq) <------> CuCl42-(aq) + CO32-(aq)

K = [CuCl42-][CO32-]/[Cl-]4 = 1.25 ×10-4M2

Initial concentration

[Cl-] = 0.650

[CuCl4-] = 0

[CO32-] = 0

change in concentration

[Cl-] = -4x

[CuCl4-] = +x

[CO32-] = +x

equillibrium concentration

[Cl-] = 0.650 - 4x

[CuCl4-] = x

[CO32-] = x

therefore

x2/(0.650 - 4x )4= 1.25×10-4

solving for x

x = 0.004468

[CO32-] = 0.004468M

Therefore

Molar solubility of CuCO3 in 0.650M NaCl solution = 0.004468M